There are 23 learners in a Grade 6 classroom; 13 boys and 10 girls. The teacher chooses two different learners at random to answer questions. Calculate the probability that both learners are girls


1
Expert's answer
2022-03-07T05:11:03-0500
"P(2\ girls)=\dfrac{\dbinom{13}{0}\dbinom{10}{2}}{\dbinom{23}{2}}=\dfrac{1(45)}{253}=\dfrac{45}{253}"